Graph the equation using the slope and the y- intercept. What is the slope and y-intercept?
y = 6/5x + 6

Answer by solver91311(24713)



In the first place, use parentheses so that there is no ambiguity in your expressions. In this case, your equation can be interpreted in three ways:



which could have been unambiguously rendered as y = (6/(5x)) + 6, or



which could have been unambiguously rendered as y = 6/(5x + 6), or



which could have been unambiguously rendered as y = (6/5)x + 6

I'm going to assume the latter because the first two are hyperbolas and do not have constant slopes.

Whenever you have an equation of the form

The slope is the coefficient on , namely the , and the -intercept is the ordered pair

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism